23. A or needs a truck for his project. Two suppliers have furnished the following data: Alternative First Cost Life Salvage Value Annual Expenses P1,200,000 P2,000,000 4 years P100,000 P180,000 P180,000 P120,000 8 years If the contractor requires a minimum retum of 18%, by how much in pesos will he be able to save upon choosing the better ltemative? $2,629,741.34 $8,167.81 $221,361.73 $372,730.03
Answer #1

If the contractor chooses Alternative A:

Initial Cost = 1200,000

Annual Expense = 180,000

Since life of the truck is 4 years, the total expense on the truck = 180,000 x 4 = 720,000

Salvage value = 100,000

Therefore, the total cost of the truck on alternative A = (Initial cost + Annual Cost) – Salvage Value

                                                                                                = (1,200,000 + 720,000) – 100,000

                                                                                                = 1,820,000

Since the Contractor makes 18% profit, therefore his savings = 18% of 1,820,000

                                                                                                               = 327,600

If the contractor chooses Alternative B:

Initial Cost = 2000,000

Annual Expense = 120,000

Since life of the truck is 8 years, the total expense on the truck = 120,000 x 8 = 960,000

Salvage value = 180,000

Therefore, the total cost of the truck on alternative A = (Initial cost + Annual Cost) – Salvage Value

                                                                                                = (2000,000 + 960,000) – 180,000

                                                                                                = 2,780,000

Since the Contractor makes 18% profit, therefore his savings = 18% of 2,780,000

                                                                                                               = 500,400

Therefore, his savings, if the contractor chooses the Alternative B = 500,400 - 327,600

                                                                                                                     = 172,800

Now, if we convert Peso 172,800 in to Dollars, it will be $8167.81. Therefore, the correct answer is $8167.81.